OpenStudy (anonymous):

I would like to know if there is an easy way to solve this or at least with th calculator......Jijiji!! ∫[4,1,(t^(7)-t^(3))/(t^(5))dx] Thanks!! (I made a mistake.)

OpenStudy (anonymous):

I gives it me another answer.. I am confused.. Please help!!

OpenStudy (anonymous):

The answer that it gave me is deifferent. Is 81/4

OpenStudy (anonymous):

I wrote it wrong... \[\int\limits_1^4 \frac{t^7-t^3}{t^5}dt=\int\limits_1^4 t^2-t^{-2} dt = \frac{t^3}{3}+t^{-1}|_1^4=\frac{1}{3}(4^3-1)+(\frac{1}{4}-1)=81/4\] http://www.wolframalpha.com/input/?i=%281%2F3%29%284^3-1%29%2B%28%281%2F4%29-1%29

OpenStudy (anonymous):

Got it, Thank you!!